OpenStudy (anonymous):

Solve |x + 6| = 6 would x= (0,1)?

myininaya (myininaya):

\[|x+6|=6 => x+6=\pm 6 \]

myininaya (myininaya):

\[x+6=6 \text{ or } x+6=-6\]

OpenStudy (maheshmeghwal9):

\[\color{red}{|x+6|=6.}\]There are 2 conditions: -\[\color{blue}{x+6=6.}\]OR\[\color{blue}{x+6=-6.}\]

myininaya (myininaya):

you must solve both of those

OpenStudy (maheshmeghwal9):

Solving these two we get?

myininaya (myininaya):

Can you solve those two equations?

OpenStudy (anonymous):

no?

myininaya (myininaya):

Lets start with this one x+6=6 We want to get x by itself but it is being added by 6 To undo addition you subtract To undo addition by 6 you subtract by 6 Whatever you do to one side you do to the other Do you think you can subtract 6 on both sides?

OpenStudy (anonymous):

yeah

myininaya (myininaya):

cool! :)

OpenStudy (maheshmeghwal9):

looks very good solution now @myininaya